OpenStudy (superhelp101):

How many liters of a 2.25 molar hydrobromic acid (HBr) solution would be needed to react completely with 100.0 grams of calcium metal? Ca (s) + 2HBr (aq) yields CaBr2 (aq) + H2 (g) 1.11 liters 2.22 liters 2.81 liters 225 liters

OpenStudy (superhelp101):

@Abhisar

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass(g/mol) Number of moles of Ca(s)=100/40.1=2.49 Mole ratio 1 mole Ca(s) reacts with 2 mole of HBr 2.49 mole of Ca(s) reacts with X mole of HBr X=2.49*2/1=4.98 Number of Moles=Concentration*Volume Volume=Number of Moles/Concentration Volume=4.98/2.25=2.216 Liters

OpenStudy (superhelp101):

Thank you! soo much!

OpenStudy (abmon98):

Your welcome :)